OpenStudy (anonymous):

find the linearization L(x) of F(x) at x=a F(x)=e^x+sin(x) when a=0

OpenStudy (anonymous):

this problem means "find the equation of the line tangent to the graph of \[y=e^x+\sin(x)\] at the point \[(0,1)\]

OpenStudy (anonymous):

you need a point and a slope. the point you already have, it is (0,1) the slope you find by taking the derivative and then replacing x by 0 to find the slope at that point

OpenStudy (anonymous):

\[f'(x)=e^x+\cos(x)\] \[f'(0)=e^0+\cos(0)=2\] so the slope is 2 and the point is (0,1) use the point slope formula to find the equation of the line

OpenStudy (anonymous):

so then you would get (y-1)=2(x-0) which would simplify to y=2x-1 which would be the same thing as L(x)=2x-1 which is the answer that is in the back of the book. thank you!!